2 ∫ 0 (x 2 + 1) dx

20 (x2 + 1) dx

Grade:12

1 Answers

Anusha
9 years ago
Ans:- 20 (x2 +1)dx
=[(x2+1 /2+1)]02
=[(x3 /3)+x]02
=(8/3)+2
=(8+6)/3
=14/3